### v2

@@ I Introduction
II Dynamically rewired 2-D small world networks
@@
In real life ecosystems with annual reproduction and dispersal, it is reasonable to assume that the majority of a species’ migration happens to neighboring habitat patches, but some individuals occasionally venture on a longer journey to distant patches. Examples span a wide variety of systems, for example trees whose seeds may be carried by the wind, aquatic larvae that may be carried by water currents, or insects that inhabit trees and may fly to distant trees. To describe this behavior, we replace a fraction $p$ of the edges in a 2-dimensional square lattice (with only nearest neighbor connections) by random connections, in a process similar to the one proposed by Watts and Strogatz [12] and that has been used to study both Ising spins [13] and other types of oscillators [14]. Note, however, that such models use a stationary network, while ours introduces different long-range connections at each iteration as will be explained next.
@@
The rewiring process is inspired by ecological dispersal and is distinct from the physically motivated quenched and annealed cases of small-world networks. Rewiring starts from a regular 2-dimensional square lattice, such that each iteration of the system’s dynamics is performed with a different network generated with the same rewiring probability $p$. In more colloquial terms, $p$ is a measure of the amount of long-range interactions, but these connections do not always happen between the same two pairs of nodes. In ecological terms, this means that individuals migrating to a distant location choose that location at random, which is a reasonable assumption given that such long-range dispersal often happens due to changing air or water currents, or simply due to diffusive motion of individuals. Similarly, nodes that are physically nearest neighbors will be connected most of the time (for small $p$), but not all the time. Rewiring the network at the beginning of every iteration, is a key difference between our model and most literature examples, which usually use quenched networks and need to perform several simulations to average over the ensemble of possible different networks [13, 14, 11]. Our rewiring process is also distinct from an annealed small world network, since rewiring is done disregarding the current state of the nodes to be rewired (see section III.2).
@@ III.1 Ecological oscillators
and the expected values are taken from the distribution of $m_{t}$ in the stationary state.
@@ IV Results
Figure 4: Binder cumulant plots for ecological oscillators (top row) and Ising spins (bottom row) for different values of rewiring probability $p$. Error bars are small and were calculated using the bootstrap method described in section 3.4.3 in [28]. The solid black triangle indicates the crossing point used to determine the critical control parameter. When the curves did not all cross at the same point, the crossing of the curves corresponding to the two largest systems was used (see the $p=0.02$ plot). The Binder cumulant values in the ordered and disordered phases are the same for the ecological oscillators as for Ising spins: $U=2/3$ in the ordered phase, and $U=0$ in the disordered phase [26].
The phase diagrams showing the separation between the ordered and disordered phases are in Fig. 5. On these graphs, the vertical axis displays the control parameter divided by its critical value for a regular lattice ($p=0$), so a direct comparison between the two different systems is straightforward. The results obtained for the regular lattice were $T=2.265\pm 0.002$ for Ising spins, consistent with the exact solution, and $\sigma=0.1483\pm 0.0009$ for ecological oscillators, consistent with the literature [6, 25]. Figure 6 shows, on log-log plots, the change in the critical control parameter for ecological oscillators and Ising spins with $p$ relative to the $p=0$ case. The vertical axis in these plots is given by
@@
Figure 5: The points indicate the critical control parameter for each value of $p$. For ecological oscillators (top), the critical noise increases monotonically as we increase $p$, with decreasing slope. For Ising spins (bottom), the critical temperature increases more rapidly, reaching a maximum value at $p=0.2$ and then decreasing slightly for higher values of $p$. The reason for this decrease is discussed in the next section. Error bars were propagated from the uncertainties in the Binder cumulant measures (Fig. 4) following the process in section 3.11 in [29]. The error bars for ecological oscillators are smaller than the data points.
Figure 6: The change in the critical control parameter (Eq. (9)) increases with $p$ according to a power law for low values of $p$. For ecological oscillators (top), the exponent of such power law is $0.510\pm 0.005$, while for Ising spins (bottom) it is $0.46\pm 0.02$. The dashed lines show the best power law fit using the four points with $p\in[0.02,0.08]$ for ecological oscillators, and the three points with $p\in[0.02,0.06]$ for Ising spins. The power law fit and uncertainties were calculated following the process in section 15.2 in [30].
@@
Now that we know how the critical point shifts with $p$, we can proceed to analyze the effects of rewiring on the order parameter. A first look at Fig. 9 shows us the intuitive result that long-range coupling stabilizes the ordered phase against noise and thermal fluctuations. However, due to the increase in the critical control parameter caused by rewiring, shown in Figs. 5 and 6, it is useful to study the order parameter in terms of the reduced control parameter, which measures how far away we are from the critical point for each value of rewiring probability. Using $f$ to represent either $\sigma$ or $T$, the reduced control parameter, $f_{r}$, is given by
@@
Figure 9: Stationary order parameter for ecological oscillators (top row) and Ising spins (bottom row) as a function of the absolute control parameter (left column) and the reduced control parameter (right column). Error bars are smaller than the symbols.
@@
Using the same reduced noise, we follow the procedure described in [35] to calculate the exponential autocorrelation time of the order parameter in the stationary state, $\tau$. Results are shown in Fig. 11, where we see that $p=0$ has an autocorrelation time one order of magnitude greater than $p=0.02$ close to criticality for both systems. If $p$ increases to 1, there is a decrease of one more order of magnitude. Furthermore, Ising spins have an autocorrelation time one order of magnitude smaller than ecological oscillators for all values of $p$.
Figure 11: Autocorrelation time of the order parameter for ecological oscillators (left) and Ising spins (right). The bottom plots are just rescaled on the vertical axis. The autocorrelation time of ecological oscillators is roughly 20 times greater than the autocorrelation time of Ising spins for the same $p$.
@@
Figure 12: Ecological oscillators have a power law with exponent $-0.63\pm 0.02$. Ising spins do not have a clear power law decay; the line shows a power law with exponent $-0.51\pm 0.02$ which is the best fit for the first three points.
@@
$G(r_{i}-r_{j})=\dfrac{1}{T}\dfrac{1}{N^{*}}\sum_{t=t_{0}}^{t_{0}+T}\sum_{i,j}x_{i,t}x_{j,t},$ (13)
where $t_{0}$ is the burning time until the stationary state is reached, $N^{*}$ is the number of terms in the inner summation, and $x$ can refer to either the two-cycle variable (Eq. (3)) or Ising spin. The value of $|r_{i}-r_{j}|$ was varied from 00 to $L/2$ considering both horizontal and vertical correlations.
@@
Figure 14: Two-point autocorrelation function with $L=128$ calculated using vertical and horizontal distances of up to $64$ lattice sites. Results for ecological oscillators are shown on top and for Ising spins at the bottom. The solid black line shows the best power law fit for short distances, with exponent $-0.257\pm 0.004$ for ecological oscillators and $-0.264\pm 0.008$ for Ising spins. The dashed line shows a power law with exponent $-0.25$ for reference. The different values of the correlation function at long distances are an effect of the finite size of the system, since $\langle m\rangle$ is not exactly zero at criticality. The decay observed at large distances is an effect of considering only horizontal and vertical correlations.
From the correlation function, we follow the method in [36] to calculate the correlation length of the system, $\xi$. Fig. 15 shows that the correlation length at criticality decays as a power law for both systems. We note, however, that for $p$ as low as $0.1$, there was no significant dependence of the correlation length with noise, with points being scattered around an average value that decreases with increasing $p$. For $p<0.1$, the correlation length peaks at the critical point, the amplitude of the peak decreases and the curve becomes noisier as we approach $p=0.1$. We also see an important effect of system size which is not captured by finite size scaling. For intermediate values of $p$ ($0.02\leq p\leq 0.2$), we see that, for smaller system sizes ($L=16$ and $32$), the correlation length is a significantly larger fraction of $L$ than for bigger systems ($L=64$ and $128$). This is simply an effect of larger systems having their effective size more drastically reduced by rewiring.
@@ V.1 Role of Memory
As noted in [25], ecological oscillators tend to maintain their phase of oscillation, which can be understood in terms of Eq. (1). As a result, ecological oscillators have a longer autocorrelation time than simple Ising spins [37, 25]. A quantitative measure of the memory present in the ecological system can be made following the work from Nareddy et al. in [37, 25]. In [37], Nareddy and Machta studied Ising spins with a self-interaction (or memory) term $K$, where a spin $s_{j}$ flips at time $t$ according not only to its nearest neighbors, but also according to its current state, described by the transition probability  [^1] Note the similarity between this model and the traditional Ising model without memory, which can be recovered if $K=0$.
@@
In the same work, they use the finite-size scaling of the critical autocorrelation time, $\tau=A(K)L^{z}$ to measure the amplitude of the divergence of the autocorrelation time as a function of memory, finding $A(K)=0.09(1+0.64e^{2K})$ (see Fig. 4 in [37]). From our Fig. 11, we see that the critical autocorrelation times of ecological oscillators are roughly $20$ times greater than those for Ising spins for most values of $p$. Since our Ising spins have $K=0$, we can use that result to infer that the ecological oscillators have $K\approx 1.9$. This is in almost perfect agreement with the memory value inferred by Nareddy et al. for ecological systems (see the rightmost point, which corresponds to the equivalent dispersal rate we use here, in figure 5b in [25]). For higher values of $p$, the critical autocorrelation time of ecological oscillators decreases more than that of Ising spins. Naturally, the explanation for this follows the same reasoning as the change in critical temperature, with Ising spins being more sensitive to the introduction on long-range interactions. From a memory perspective, this indicates that larger values of $p$ lead to a weaker effect of memory in ecological oscillators. For $p=1$, when the critical autocorrelation time of ecological oscillators is about 10 times that of Ising spins, the value of K would be $K\approx 1.6$. This discrepancy is likely due to the higher values of control parameter for greater values of $p$.
@@
In Fig. 5, we see that both systems display an increase in the critical control parameter as we increase the fraction of long-range coupling, which is the expected effect of rewiring, since it facilitates long-range ordering. A deeper look into those results reveals another display of the memory of ecological oscillators, considering that Ising spins respond more strongly to a small increase in the fraction of long-range interactions. This can be explained by the fact that the dynamics of ecological oscillators depends not only on the nodes to which they are currently coupled, which directly determine the population via Eq. (2), but also on nodes to which they were previously coupled, which played a role in determining the population at previous time steps and is thus encoded in the oscillator’s memory. Averaging a node’s connections over time for small $p$ will show us that a node is mostly connected to its nearest neighbors, such that if the system has some memory of its past, it will not respond to rare long-range interactions as dramatically as a system without memory. At high values of $p$, we see another consequence of memory: the critical control parameter remains increasing for ecological oscillators, while it decreases for the Ising model (for $p>0.2$). This effect is perhaps caused by the stochasticity introduced by the rewiring process, since it drives the system out of equilibrium as shown in Fig. 3, requiring lower temperatures to achieve the ordered state. For ecological oscillators, the critical noise level increases monotonically with $p$, with decreasing slope, because the system has memory, so even with new neighbors at each time step, it is easier to synchronize the network because nodes retain information from their previous connections. This means that the equilibrium state of the ecological system is not as disturbed by rewiring, and is therefore closer to the standard fully connected mean field model. A final notable feature shown in Fig. 5 is that the critical control parameter increases by roughly $25\%$ at its maximum value for both systems. This is significantly less than the increase observed in quenched networks [13], and the discrepancy can be understood in terms of the added stochasticity due to the rewiring dynamics.
@@
It is also interesting to notice that the power laws observed in Fig. 6 are compatible with an exponent of $1/2$, which was also observed for Ising spins on quenched small-world networks in [13], for which the exponent was found to be $0.52\pm 0.03$. The similarity emerges despite the different network dynamics, both in terms of local dynamics (ecological oscillators vs Ising spins) and network structure (dynamic vs static rewiring). A satisfactory explanation for such exponent will be the subject of future studies. Similar results have been observed for different systems in different dimensions. Namely, a logarithmic increase in the critical temperature of the XY model was observed when rewiring a 1-D ring lattice [11], and a power law increase in critical temperature, with exponent $0.96\pm 0.04$, was observed for the Ising model when rewiring a 3-D lattice [13]. This allows us to conjecture that the power law may be described by an exponent given by $(d-1)/2$, where $d$ is the dimensionality of the lattice being rewired.

### v3

@@ I Introduction
II Dynamically rewired 2-D small-world networks
@@
In real life ecosystems with annual reproduction and dispersal, it is reasonable to assume that the majority of a species’ migration happens to neighboring habitat patches, but some individuals occasionally venture on a longer journey to distant patches. Examples span a wide variety of systems, such as trees whose seeds may be carried by the wind, aquatic larvae that may be carried by water currents, or insects that inhabit trees and may fly to distant trees. To describe this behavior, we replace a fraction $p$ of the edges in a 2-dimensional square lattice (with only nearest neighbor connections) by random connections, in a process similar to the one proposed by Watts and Strogatz [12] and that has been used to study both Ising spins [13] and other types of oscillators [14]. Note, however, that such models use a stationary network, while ours starts with a regular underlying structure that at every iteration is rewired, introducing long-range connections that only exist at each iteration of the system’s dynamics.
@@
The rewiring process is inspired by ecological dispersal and is distinct from the physically motivated quenched and annealed cases of small-world networks. Rewiring always starts from a regular 2-dimensional square lattice, such that each iteration of the system’s dynamics is performed with a different network generated with the same rewiring probability $p$. In more colloquial terms, $p$ is a measure of the amount of long-range interactions, but these connections do not always happen between the same two pairs of nodes. In ecological terms, this means that individuals migrating to a distant location choose that location at random, which is a reasonable assumption given that such long-range dispersal often happens due to changing air or water currents, or simply due to diffusive motion of individuals. Similarly, nodes that are physically nearest neighbors will be connected most of the time (for small $p$), but not all the time. Rewiring the network at the beginning of every iteration, is a key difference between our model and most literature examples, which usually use quenched networks and need to perform several simulations to average over the ensemble of possible different networks [13, 14, 11]. Our rewiring process is also distinct from an annealed small-world network, since rewiring is done disregarding the current state of the nodes to be rewired (see section III.2).
@@ III.1 Ecological oscillators
and the expected values are taken from the distribution of $m_{t}$ in the stationary state. The crossing point of Binder cumulant curves for different system sizes are known to be a highly reliable estimator of the critical point [26].
@@ IV Results
Figure 4: Binder cumulant plots for ecological oscillators (top row) and Ising spins (bottom row) for different values of rewiring probability $p$. Error bars are small and were calculated using the bootstrap method. This method consists of resampling the data with replacement, averaging them, and taking the standard deviation of the averages after a suitably large number of resamples [28]. The solid black triangle indicates the crossing point used to determine the critical control parameter. When the curves did not all cross at the same point, the crossing of the curves corresponding to the two largest systems was used (see the $p=0.02$ plot). The Binder cumulant values in the ordered and disordered phases are the same for the ecological oscillators as for Ising spins: $U=2/3$ in the ordered phase, and $U=0$ in the disordered phase [26].
The phase diagrams showing the separation between the ordered and disordered phases are in Fig. 5. On these graphs, the vertical axis displays the control parameter divided by its critical value for a regular lattice ($p=0$), so a direct comparison between the two different systems is straightforward. The critical control parameters obtained for the regular lattice were $T_{c}(p=0)=2.265\pm 0.002$ for Ising spins, consistent with the exact solution, and $\sigma_{c}(p=0)=0.1483\pm 0.0009$ for ecological oscillators, consistent with the literature [6, 25]. The bottom panels in Fig. 5 show, on log-log plots, the change in the critical control parameter for ecological oscillators and Ising spins with $p$ relative to the $p=0$ case. The vertical axis in these plots is given by
@@
Figure 5: The points in the top plots indicate the critical control parameter for each value of $p$. For ecological oscillators (left), the critical noise increases monotonically as we increase $p$, with decreasing slope. For Ising spins (right), the critical temperature increases more rapidly, reaching a maximum value at $p=0.2$ and then decreasing slightly for higher values of $p$. The reason for this decrease is discussed in the next section. Error bars were propagated from the uncertainties in the Binder cumulant measures (Fig. 4) following the process in [29]. The error bars for ecological oscillators are smaller than the data points. The bottom plots show the change in the critical control parameter (Eq. (9)), which increases with $p$ according to a power law for low values of $p$. For ecological oscillators, the exponent of such power law is $0.510\pm 0.005$, while for Ising spins it is $0.46\pm 0.02$. The dashed lines show the best power law fit using the four points with $p\in[0.02,0.08]$ for ecological oscillators, and the three points with $p\in[0.02,0.06]$ for Ising spins. The power law fit and uncertainties were calculated following the process in [30].
@@
Now that we know how the critical point shifts with $p$, we can proceed to analyze the effects of rewiring on the order parameter. Due to the increase in the critical control parameter caused by rewiring, shown in Fig. 5, it is useful to study the order parameter in terms of the reduced control parameter, which measures how far away we are from the critical point for each value of rewiring probability. Using $f$ to represent either $\sigma$ or $T$, the reduced control parameter, $f_{r}$, is given by
@@
Using the same reduced noise, we calculated the integrated autocorrelation time of the order parameter in the stationary state, $\tau=\sum_{t=0}^{T}\rho_{m}(t)$, where $\rho_{m}(t)$ is the normalized autocorrelation function of the order parameter [35] and $T$ is an appropriate time lag [36]  [^1] Following [35, 36], we calculated $\tau$ iteratively. Starting with $T=10$, we set $T=6\tau$ and updated the value of the autocorrelation time until it converged or we summed over the entire length of the autocorrelation function, which indicates a divergence of $\tau$. Results are shown in Fig. 9, where we see that $p=0$ has an autocorrelation time one order of magnitude greater than $p=0.02$ close to criticality for both systems. If $p$ increases to 1, there is a decrease of one more order of magnitude. Furthermore, Ising spins have an autocorrelation time one order of magnitude smaller than ecological oscillators for all values of $p$.
Figure 9: Autocorrelation time of the order parameter for ecological oscillators (left) and Ising spins (right). The bottom plots have a rescaled vertical axis and correspond to the shaded region in the top plots. The autocorrelation time of ecological oscillators is roughly 20 times greater than the autocorrelation time of Ising spins for the same $p$.
@@
Figure 10: Ecological oscillators have power law with exponent $-0.63\pm 0.02$. Ising spins do not have a clear power law decay; the line shows a power law with exponent $-0.51\pm 0.02$ which is the best fit for the first three points.
@@
$G(|\mathbf{r_{i}-r_{j}}|)=\dfrac{1}{T}\dfrac{1}{N^{*}}\sum_{t=t_{0}}^{t_{0}+T}\sum_{i,j}x_{i,t}x_{j,t},$ (13)
where $t_{0}$ is the burning time until the stationary state is reached, $N^{*}$ is the number of terms in the inner summation, and $x$ can refer to either the two-cycle variable (Eq. (3)) or Ising spin. The vector $\mathbf{r_{i}}$ is the position vector or site $i$ and the value of $\mathbf{r_{i}-r_{j}}$ was varied from 00 to $L/2$ and taken only in the vertical and horizontal directions.
@@
Figure 12: Critical two-point correlation function with $L=128$. The correlation function is measured using only vertical and horizontal spatial lags of up to $64$ lattice sites. Results for ecological oscillators are shown on top and for Ising spins at the bottom. The solid black line shows the best power law fit for short distances, with exponent $-0.257\pm 0.004$ for ecological oscillators and $-0.264\pm 0.008$ for Ising spins. The dashed line shows a power law with exponent $-0.25$ for reference. The different values of the correlation function at long distances are an effect of the finite size of the system, since $\langle m\rangle$ is not exactly zero at criticality. The decay observed at large distances is an effect of considering only horizontal and vertical correlations.
The correlation function decays exponentially at large distances with a length scale of the correlation length, $\xi$. To determine $\xi$, we first calculate the Fourier transform of the correlation function and fit the points with low wave number to a Lorentzian function, whose coefficients can be combined to obtain $\xi$ [38, 36]. Fig. 13 shows that the correlation length at criticality decays as a power law for both systems. We note, however, that for $p$ as low as $0.1$, there was no significant dependence of the correlation length with noise, with points being scattered around an average value that decreases with increasing $p$. For $p<0.1$, the correlation length peaks at the critical point, the amplitude of the peak decreases and the curve becomes noisier as we approach $p=0.1$. We also see an important effect of system size which is not captured by finite size scaling. For intermediate values of $p$ ($0.02\leq p\leq 0.2$), we see that, for smaller system sizes ($L=16$ and $32$), the correlation length is a significantly larger fraction of $L$ than for bigger systems ($L=64$ and $128$). This is simply an effect of larger systems having their effective size more drastically reduced by rewiring.
@@ V.1 Role of Memory
As noted in [25], ecological oscillators tend to maintain their phase of oscillation, which can be understood in terms of Eq. (1). As a result, ecological oscillators have a longer autocorrelation time than simple Ising spins [39, 25]. Although the dynamics of ecological oscillators is Markovian, there is an effective memory of the phase of oscillation that is not present in the standard dynamics of Ising spins. A quantitative measure of the memory present in the ecological system can be made following the work from Nareddy et al. in [39, 25]. In [39], Nareddy and Machta studied Ising spins with a self-interaction (or memory) term $K$, where a spin $s_{j}$ flips at time $t$ according not only to its nearest neighbors, but also according to its current state, described by the transition probability  [^2] Note the similarity between this model and the traditional Ising model without memory, which can be recovered if $K=0$.
@@
In the same work, they use the finite-size scaling of the critical autocorrelation time, $\tau=A(K)L^{z}$ to measure the amplitude of the divergence of the autocorrelation time as a function of memory, finding $A(K)=0.09(1+0.64e^{2K})$ (see Fig. 4 in [39]). From our Fig. 9, we see that the critical autocorrelation times of ecological oscillators are roughly $20$ times greater than those for Ising spins for most values of $p$. Since our Ising spins have $K=0$, we can use that result to infer that the ecological oscillators have $K\approx 1.9$. This is in almost perfect agreement with the memory value inferred by Nareddy et al. for ecological systems (see the rightmost point, which corresponds to the equivalent dispersal rate we use here, in figure 5b in [25]). For higher values of $p$, the critical autocorrelation time of ecological oscillators decreases more than that of Ising spins. Naturally, the explanation for this follows the same reasoning as the behavior of the critical temperature, with Ising spins being less sensitive to an increase in the fraction of long range interactions once $p$ reaches a threshold close to $0.2$. From a memory perspective, this indicates that larger values of $p$ lead to a weaker effect of memory in ecological oscillators. For $p=1$, when the critical autocorrelation time of ecological oscillators is about 10 times that of Ising spins, the value of K would be $K\approx 1.6$. This discrepancy is likely due to the higher values of control parameter for greater values of $p$.
@@
In Fig. 5, we see that both systems display an increase in the critical control parameter as we increase the fraction of long-range coupling, which is the expected facilitation of long-range ordering due to rewiring. A deeper look into those results reveals another display of the memory of ecological oscillators, considering that Ising spins respond more strongly to a small increase in the fraction of long-range interactions. This can be explained by the fact that the dynamics of ecological oscillators depends not only on the nodes to which they are currently coupled, which directly determine the population via Eq. (2), but also on nodes to which they were previously coupled, which played a role in determining the population at previous time steps and is thus encoded in the oscillator’s memory. Averaging a node’s connections over time for small $p$ will show us that a node is mostly connected to its nearest neighbors, such that if the system has some memory of its past, it will not respond to rare long-range interactions as dramatically as a system without memory. At high values of $p$, we see another consequence of memory: the critical control parameter remains increasing for ecological oscillators, while it decreases for the Ising model (for $p>0.2$). This effect is perhaps caused by the stochasticity introduced by the rewiring process, since it drives the system out of equilibrium as shown in Fig. 3, requiring lower temperatures to achieve the ordered state. For ecological oscillators, the critical noise level increases monotonically with $p$, with decreasing slope, because the system has memory, so even with new neighbors at each time step, it is easier to synchronize the network because nodes retain information from their previous connections. This means that the equilibrium state of the ecological system is not as disturbed by rewiring, and is therefore closer to the standard fully connected mean field model. It would be interesting in future work to carry out a numerical study of the Ising model with memory on rewired networks to confirm that it behaves similarly to ecological oscillators. A final notable feature shown in Fig. 5 is that the critical control parameter increases by roughly $25\%$ at its maximum value for both systems. This is significantly less than the increase observed in quenched networks [13], and the discrepancy can be understood in terms of the added stochasticity due to the rewiring dynamics.
@@
It is also interesting to notice that the power laws observed in the bottom panels of Fig. 5 are compatible with an exponent of $1/2$, which has been previously observed in static small-world networks to be $0.52\pm 0.03$ [13], although the range of $\log p$ is too small to confirm a true power law. Similar caveats apply to the power laws seen in Figs. 10 and 13. The similarity emerges despite the different network dynamics, both in terms of local dynamics (ecological oscillators vs Ising spins) and network structure (dynamic vs static rewiring). A satisfactory explanation for such exponent will be the subject of future studies. Similar results have been observed for different systems in different dimensions. Namely, a logarithmic increase in the critical temperature of the XY model was observed when rewiring a 1-D ring lattice [11], and a power law increase in critical temperature, with exponent $0.96\pm 0.04$, was observed for the Ising model when rewiring a 3-D lattice [13]. This allows us to conjecture that the power law may be described by an exponent given by $(d-1)/2$, where $d$ is the dimensionality of the lattice being rewired.